OpenStudy (christos):

Tanθ=2. 180*<θ<270*. Cosθ=?

hartnn (hartnn):

use, \(\sec^2 \theta =1+\tan^2 \theta \) to find sec theta first.

hartnn (hartnn):

what u get ?

OpenStudy (christos):

Same form apply if eg i have csc and cot?

hartnn (hartnn):

cot is easy cot x = 1/ tan x for csc , you first find cot and then use \(\csc^2x = 1+\cot^2x\)

hartnn (hartnn):

for your initial proble, once you find sec theta, then cos theta = 1/sec theta and since in 3rd quadrant cos is negative, add a - *minus* sign to it

OpenStudy (christos):

Tanθ= 2 180<θ<270. Cosθ=?

hartnn (hartnn):

yeah, i know the Q, its posted above also :)

hartnn (hartnn):

did u get the value of sec theta ??

OpenStudy (christos):

On which quadrant is sec(θ)?

hartnn (hartnn):

sec θ is just a ratio , it cannot be in any quadrant, θ is an angle, it can be in a quadrant 180*<θ<270* implies its in 3rd quadrant.

hartnn (hartnn):